The speed of light in air is 3.0x10^8ms/2. Calculate the refractive index of the material which the fibre is made.


1
Expert's answer
2021-05-05T13:30:12-0400

Given,

Speed of light in air "v_1=3\times 10^8\ m/s"

Let speed of light in fibre is "v_2=1.8\times 10^8\ m/s"

So,

Refractive index of fibre "n_g=\dfrac{speed \ of light\ in\ vacuum}{speed\ of \ light\ in\ fibre}"

"n_g=\dfrac{3\times 10^8}{1.8\times 10^8}=1.67"

So, Refractive Index of fibre is 1.67